### v1

# Scene: Procedural Controllable Large-scale Scene Generation via Large-language Models

Abstract
Due to its great application potential, large-scale scene generation has drawn extensive attention in academia and industry. Recent research employs powerful generative models to create desired scenes and achieves promising results. However, most of these methods represent the scene using 3D primitives (e.g. point cloud or radiance field) incompatible with the industrial pipeline, which leads to a substantial gap between academic research and industrial deployment. Procedural Controllable Generation (PCG) is an efficient technique for creating scalable and high-quality assets, but it is unfriendly for ordinary users as it demands profound domain expertise. To address these issues, we resort to using the large language model (LLM) to drive the procedural modeling. In this paper, we introduce a large-scale scene generation framework, Scene, which can automatically produce high-quality procedural models according to designers’ textual descriptions.Specifically, the proposed method comprises two components, PCGBench and PCGPlanner. The former encompasses an extensive collection of accessible procedural assets and thousands of hand-craft API documents. The latter aims to generate executable actions for Blender to produce controllable and precise 3D assets guided by the user’s instructions. Our Scene can generate a city spanning 2.5 km 2.5 km with delicate layout and geometric structures, drastically reducing the time cost from several weeks for professional PCG engineers to just a few hours for an ordinary user. Extensive experiments demonstrated the capability of our method in controllable large-scale scene generation and editing, including asset placement and season translation. Our project page: https://scenex-lab.github.io/.
Keywords:
Large-scale Scene Generation LLM Agents PCG
1 Introduction
With the rapid progress in generative modeling, 3D content generation has witnessed remarkable achievements in automatic creation of 3D objects [39, 43, 35, 36], avatars [58, 27, 17], and scenes [31, 52, 29, 9]. In this work, we focus on automatic 3D scene generation, which has extensive applications in domains like gaming, animation, film, and virtual reality. Although previous approaches have achieved promising results, a significant gap still remains for industrial-grade applications due to incompatible representations with the industrial pipeline of the scene using 3D primitives (e.g. point cloud or radiance field).
To bridge the gap, we resort to procedural modeling, an advanced technique for the fast creation of high-quality 3D assets. It can generate realistic and intricate 3D assets in natural scenes through adjustable parameters and rule-based systems [32] using software like Blender. For example, Infinigen [42] proposes a procedural generator that combines different procedural modeling algorithms to generate large-scale natural scenes encompassing terrain, weather, vegetation, and wildlife. [33], [5] and [46] use procedural modeling to generate city-level street or layout. These procedural approaches generate high-quality 3D assets at the object, scene, and city levels. However, the comprehensive grasp of generation rules, algorithmic frameworks, and individual parameters of procedural modeling make the generation process beginner-unfriendly and time-consuming, especially for large-scale scene or city generation. For instance, generating a New York city, as shown in Fig 1, requires the effort of a professional PCG engineer working for over two weeks. Moreover, due to the substantial learning curve for users, it is formidable to generate large-scale scenes using procedural modeling.
To address the above problems, existing works such as 3D-GPT [45] and SceneCraft [19] introduce an instruction-driven 3D modeling method by integrating an LLM agent with procedural generation software. Despite the successful collaboration with human designers to establish a procedural generation framework, these methods still exhibit significant limitations. Firstly, SceneCraft is unable to achieve fine-grained 3D asset editing due to the uneditable nature of fixed 3D assets compared to those created by procedural generation. Secondly, 3D-GPT, being based on the parametric procedural generation model Infinigen, restricts its ability to fully utilize existing PCG resources. Additionally, it is incapable of generating high-quality and large-scale scenes, especially expansive terrains and boundless cities. Lastly, the diversity of assets is limited due to constraints imposed by prompt expressions.
To address these issues, we propose an instructive large-scale scene generation framework, Scene, which can automatically generate high-quality scenes according to designers’ textual descriptions. By introducing a PCGBench, we collect extensive accessible procedural assets, including various terrains, cities, and natural objects like plants and animals. This helps us to generate diverse, realistic, and practical large-scale scenes. Moreover, we also construct thousands of API parameter documents, which are used to formulate the agent planner process and provide a quantitative evaluation. Inspired by existing LLM-driven planning methods [49, 44, 51, 28], we develop a PCGPlanner framework to produce correct API actions interfaced with Blender, to automatically generate controllable and precise 3D assets guided by users’ instructions. This LLM-based automated solution significantly enhances the efficiency of scene generation, reduces the learning curve for procedural scene modeling, and enables the customized creation of target scenes. In addition to the inherent advantages of procedural modeling such as photorealistic and real geometry, our Scene possesses several key properties:
1.
Efficiency: Benefiting from the proposed PCGBench and PCGPlanner, our Scene can rapidly produces extensive, high-quality 3D assets, including terrain, city, and forest. Moreover, we only need a few hours to generate a large-scale city, whereas it would take a professional designer over two weeks.
2.
Controllability: Scene can generate 3D scenes satisfying personalized demands. For example, we can achieve scene editing according to the corresponding descriptions, such as adding objects, placing objects at a location, translating the season, and so on.
3.
Diversity: Scene breaks the limitations of specific generation algorithms and predefined rules, leveraging advanced language model technology to achieve more flexible and diverse large-scale scene generation. Specifically, we generate large-scale terrains in the natural world, like deserts, lakes, forests, and mountains, and high-quality cities in the real world, such as New York, Paris, and Tokyo, as well as various objects, including plants, animals, and buildings.
2 Related Works
2.1 Learning Based 3D Generation.
In recent years, 3D asset generation has witnessed rapid progress, combining the ideas of computer graphics and computer vision to realize the free creation of 3D content. Presently, predominant research in 3D asset generation primarily concentrate on creating individual objects [39, 35, 30, 8], 3D avatars [58, 27, 18, 25], and 3D scenes [16, 61, 11, 62, 59]. Among these, ZeroShot123 [35] proposes a method based on a diffusion model, which realizes the 3D model of the target based on a picture. DreamFusion [39] proposes a NeRF-based approach that allows the model to generate a corresponding 3D model based on the input text.
Compared with the generation of a single object, the generation of large-scale scenes, including the generation of natural landforms [15, 34, 9] and borderless cities [31, 29, 52], is undoubtedly more practical and challenging in the metaverse and game production. CityDreamer [52] builds vast, large-scale 3D cities based on the layout of real cities, enhancing the accuracy and stability of urban reconstruction. SceneDreamer [9] proposes a method to generate 3D borderless scenes within 2D plots using BEV representations. These learning-based methods do not have geometric structures, making them difficult to apply directly to virtual engines such as UE. For example, although InfiniCity [31] and CityDreamer [52] can generate visually appealing urban scenes, the absence of clear geometry may lead to issues in practical applications, such as intersecting objects and spatial discontinuities.
2.2 Procedural Based 3D Generation.
Researchers have delved into the procedural generation of natural scenes [12, 60] and urban scenes [33, 46, 48, 54] using Procedural Content Generation (PCG). For instance, PMC [38] proposes a procedural way to generate cities based on 2D ocean or city boundaries. It employs mathematical algorithms to generate blocks and streets and utilizes tailgating technology to generate the geometry of buildings. While these traditional computer graphic methods can produce high-quality 3D data, all parameters must be pre-entered into the procedurally generated process. The resultant 3D data is bound by rule limitations, introducing a certain level of deviation from the real world. This significantly constrains flexibility and practical usability. Infinigen [42] introduces a technique for procedurally generating realistic 3D natural objects and scenes. This methodology facilitates the programmatic generation of all assets, with meshes and textures generated through random mathematical algorithms. Although Infinigen generates infinitely real assets, users are unable to customize the generated outcomes because of their specific requirements. Therefore, we propose a more convenient procedural generation method to produce high-quality 3D assets.
2.3 LLM Based Task Planner
Benefiting from knowledge hidden in the large-language model (LLMs) [41, 1, 10, 26, 3], researchers explore LLMs to address intricate tasks beyond canonical language processing domains. These tasks encompass areas such as mathematical reasoning [22, 50], medicine [53, 23], and planning [57, 14, 21, 20]. Thanks to powerful reasoning and generalization capabilities, LLMs act as practiced planners for different tasks. For example, [20] utilizes the expansive domain knowledge of LLMs on the internet and their emerging zero-shot planning capabilities to execute intricate task planning and reasoning. [14] investigates the application of LLMs in scenarios involving multi-agent coordination, covering a range of diverse task objectives. [56] presents a modular framework that employs structured dialogue through prompts among multiple large pretrained models. Moreover, specialized LLMs for particular applications have been explored, such as HuggingGPT [44] for vision perception tasks, VisualChatGPT [51] for multi-modality understanding, Voyager [49] and [63], SheetCopilot [28] for office software, and Codex [6] for Python code generation. Inspired by the existing works, we explore the LLM agent to PCG software, e.g., Blender, to provide automatic 3D assets generation.
3 Methods
The proposed Scene framework contains two main components: the PCGBench and PCGPlanner. PCGBench is a large-scale dataset that includes diverse resources in procedural modeling, which is used to perform as an asset bank and evaluate the generation performance. Specifically, it includes collecting plugins and assets from the Internet, asset annotations, and hand-craft API documents. Based on our PCGBench, a novel PCGPlanner framework is proposed to automatically generate 3D scenes using both PCG and pre-defined models. It comprises a task planner stage, an asset retrieval stage, and an action execution stage guided by agents with custom-built instructions. By the cooperation between the PCGBench and novel PCGPlanner agents, Scene is endowed with the capability of efficient, controllable generation of diverse, high-quality scenes.
3.1 PCGBench
To build PCGBench, we first collect abundant resources of PCG models from Blender Market111https://blendermarket.com/, which includes action APIs and 3D assets. The action APIs consist of raw APIs and detailed parameters that are used to control asset generation. Specifically, we have collected 1,532 individual raw APIs in total, involving each aspect in the scene-generation tasks, such as controlling generated assets’ placement, color, size, and height. Then, we manually merge the most essential raw APIs according to their functions and obtain 45 action functions. Each action function contains multiple raw APIs with the corresponding parameters to produce a complete 3D asset. For example, to generate a 3D tree as Fig 2 shows, we define two individual action functions such as tree_mesh_generate and tree_material_generate. Each function comprises several raw APIs with the corresponding parameters like TreeHight, CrownDensity, LeafDensity, etc. These parameters can be manipulated at the action input, and we can generate various trees by changing these parameters.
To achieve controllable 3D model generation guided by human instructions, we should formulate the action functions for large-language models (LLMs), e.g., GPT-4222https://openai.com/gpt-4, following previous works [49, 45, 6]. To achieve this, we compile lots of action functions guidance documents, denoted as API Documents, to deploy our functions for LLMs. Concretely, this document includes information on API types, function descriptions, API names, and a set of raw APIs with the corresponding parameter details. Besides, we also provide a comprehensive record of all parameter names and their possible values. The details are referred to Fig 2. Moreover, we have categorized these documents into four modules:{ Terrain, City, Detail, Weather}. Simultaneously, we have developed a concise document outlining the scene generation process, providing detailed explanations of the construction sequence and capability boundaries for each module.
Regarding asset collection, we have meticulously curated an extensive 3D dataset comprising 1,908 finely crafted models and 1,294 textures. This diverse collection spans categories including buildings, flora, urban infrastructure, and textures gathered from the vast expanse of the Internet. These existing assets provide an alternative solution for 3D scene generation by retrieving existing individual procedural models rather than progressively generation, which can significantly improve the efficiency of creating large-scale scenes. For retrieval, we use CLIP [40] encoder feature, achieving text-to-text retrieval and text-to-image retrieval. Therefore, we also provide abundant annotations for collected 3D assets by rendering a picture for each 3D asset.
3.2 PCGPlanner
To achieve LLM-driven asset generation, we introduce PCG Planner Agents’ hierarchy, which is comprised of multiple agents that can automatically generate high-quality and large-scale scenes corresponding to the users’ textual descriptions. Following previous works [44, 51, 49, 6], we design our module encompassing three primary stages: task planning, asset retrieval, and action execution. In the task planning stage, a sequential planning agent pair is introduced to process the user query and output a detailed sub-task list. Then, to precisely connect the sub-task with our PCGBench, we retrieve the actions and assets mentioned in the task list. Finally, the execution agent leverages the annotated action APIs to process subtasks of procedural generation or asset manipulation.
Agent construction To endow the LLMs with the capability for modeling scenes like a specialist, we introduce a systematic template to prompt LLMs as agents to achieve the subtasks in a scene modeling task. For each agent, the prompt has a similar structure defined as (, , , , ), where i {dispatch, specialist, retrieval, execution} distinguishes the responsibility of agents. The constituents of the prompt are defined as follows:
Role Each agent is given a specific role R which describes its responsibility in the scene generation process.
Task T gives a detailed explanation of the goals for the agent. In the meanwhile, the constraints are also expounded executing these tasks.
Document At each step, the agent is prompted by a knowledge document according to their task. We denote D as the collection that contains all the knowledge documentation pre-defined in PCGBench.
Format F denoted as the output format for each agent. To precisely and concisely convey information between agents, the output format of each agent is strictly defined.
Examples To help agents respond strictly following the format, we demonstrated several examples E for reference in each prompt.
3.2.1 Task Planner
For a scene generation task, multiple action APIs from different modules are usually needed. There is a great possibility that different APIs require to be called in a sequential order. Thus, a correct function arrangement between different modules and APIs is of essential importance. To ensure competence in handling a broad spectrum of scene-generation tasks and surmounting token limitations, we present a two-stage iterative task planner comprising a dispatch agent and a specialist agent as depicted in Fig. 3.
Dispatch agent Dispatch agent is the first agent to receive users’ descriptions of the target scene, it is responsible for extracting information from the sentences. It serves as a planner to interpret user intents into a rough plan and list all the objects mentioned in the input instruction. In the meantime, users may omit some common background objects in the description, the agent will automatically add all these objects that need to be taken into account in the rough plan.
The detailed object requirements of a scene and their module correspondence principle are described in the knowledge documentation . We formulate the dispatch agent as follows:
where is the user input, stands for an object listed in the rough plan in the formation of {Module: Object name}
Specialist agent Then, the specialist agent is activated to make further detailed plans.
At this stage, a specialist agent serves as a professional designer who is capable of giving detailed plans for the objects given in the rough plan. According to the module class of the object, the specialist agent is prompted with the corresponding knowledge documentation , where contains the generation mechanisms and explanations for each module. The specialist agent decomposes each rough plan into a detailed plan with several sub-tasks, where each subtask corresponds to an executable API or an asset to be imported
where stands for one of the sub-tasks to be completed.
3.2.2 Asset Retrieval
With the detailed plan given by the specialist agent, we need to search in our PCGBench for a suitable API to execute or an asset to import. To precisely retrieve them, we propose a retrieval agent that identifies whether the target of the sub-task type is an asset or an API. With the prompt , the agent is also instructed to output a description as the search index of the asset or the API. The description is then encoded into an embedding for the retrieval. Here, a pre-trained CLIP [40] is used to search our action APIs by text-to-text retrieval and 3D models by text-to-image retrieval, respectively. As shown in the right part of Fig 2, the rendering images of each asset in our PCGBench are encoded into normalized 768-dimensional vectors, and then compared with the input description embedding. One of the five most similar results is randomly selected as the outcome and imported into the Blender scene. The best-matched result is selected by calculating the similarity by cosine distance.
where stands for an retrieved API, stands for the API collection. is the retrieved asset from asset collection . For text-to-text retrieval, the most relevant API is also selected by its functional description embedding. If the retrieval result is an asset, it will be directly imported into the scene project. On the other hand, the retrieved API is to be passed to the action execution agent.
3.2.3 Action Execution
In the PCGBench, each action API is defined in the form of a Python function with multiple pre-defined parameters. After obtaining the , we need to make custom modifications for these initial parameters to meet the task description. For each API , there is a knowledge documentation that passed into the prompt . As illustrated in Fig 2 (a), the knowledge documentation of each action API is recorded in a JSON format and has the same architecture. Take tree_mesh_generate API as an example, each parameter in the function is precisely defined and explained in tree_mesh_doc = {"type": "tree mesh", "Description": "Function for generating trees mesh.", "Function API": "tree_mesh_generate", "Parameter Information": {"Possible values": ("High", "Medium", "Low"), … } }. The execution agent can be expressed as:
Where is the manipulated version of retrieved API . There is a chance that the descriptions contain less information for all the parameters. In this situation, we instruct the execution agent to automatically choose a feasible parameter from the options according to the other parameters. The execution agent ultimately executed the action API with the inferred parameters in Blender Python environments, incrementally adding new objects and making modifications in scene project until all the tasks were completed.
4 Experiments
The goals of our experiments are threefold: (i) to verify the capability of Scene for generating photorealistic large-scale scenes, including nature scenes and cities; (ii) to demonstrate the effectiveness of Scene for personalized editing, such as adding or changing; (iii) to compare different LLMs on the proposed benchmark.
4.1 Benchmark Protocol
Dataset. To evaluate the effectiveness of proposed Scene, we use GPT-4 to generate high-quality 50 scene descriptions, 50 asset descriptions, and 20 asset editing descriptions. The scene descriptions involve natural scenes and cities. Then, we feed them to our Scene to generate corresponding models, which are used to perform quantitative and qualitative comparisons.
Models. When generating and editing the 3D scenes, we adopt the leading GPT-4 as the large language model with its public API keys. To ensure the stability of LLM’s output, we set the decoding temperature as 0.
Metrics: We use Executability Rate (ER@1) and Success Rate (SR@1) to evaluate the capabilities of LLMs on our Scene. The former measures the proportion of proposed actions that can be executed, and the latter is used to evaluate action correctness [7]. Moreover, to quantify the aesthetic level of the generated scenes, we adopt a unified judgment standard as a reference. We divide the aesthetics of generated scenes into five standards: Poor (1-2 points)/Below Average (3-4 points)/Average (5-6 points)/Good (7-8 points)/Excellent (9-10 points). We enlisted 35 volunteers to assess the quality of our generation, including 5 PCG experts. We compute average score (AS) and average expert score (AES) to evaluate the effectiveness of our method.
4.2 Main Results
Large-Scale Scene Generation. To intuitively illustrate the capability of Scene in generating large-scale scenes, we select three subsets with significant differences for comparison. We perform quantitative and qualitative results, including generation quality, aesthetic evaluation and efficacy comparison.
Generation Quality. We first provide several examples of our Scene for generating large-scale nature scenes and unbounded cities. The results are shown in Fig 4. From the results, we can observe that the proposed Scene can produce highly realistic scenes in both natural scenes and cities. Moreover, the generated content is correctly corresponding to the provided texture descriptions. These demonstrate the power and effectiveness of our proposed LLM-driven automatic 3D scene generation framework.
Furthermore, to compare with existing learning-based methods, we provide the qualitative comparison with several representing methods on 3D city generation, such as Persistent Nature [4], SceneDreamer [9],InfiniCity[31], and CityDreamer [52]. Fig. 5 shows the qualitative comparison results between the learning-based methods for city scene reconstruction work and Scene. From the results, we can observe that learning-based methods commonly suffer similar problems: low 3D consistency and building structural distortions. For example, PersistanNature [4] and InfiniCity [31] both appear to have severe deformation in the whole scene level. SceneDreamer [9] and CityDreamer [52] have better structure consistency, but the building quality is still relatively low. These factors limit their large-scale application in industry. In comparison, the proposed Scene generates highly realistic and well-structured urban scenes without the issues of structural distortions and layout defects compared to learning-based methods. These results demonstrate the effectiveness of our method for large-scale city scene generation.
Aesthetic Evaluation. To better evaluate the generation quality of Scene, we collect the results of related works involving text-to-3D work and Blender-driven 3D generation. These results are subjected to aesthetic evaluation by a panel comprising of 35 voluntary contributors and 5 experts in 3D modeling, with the scoring criteria outlined in Section 4.1. As shown in Table 2, our scores for AS and AES surpass the second-highest scores by 2.10 and 0.9 points, respectively. Compared to the average level of other works, our project reaches a good level, indicating the high generation quality of Scene.
To evaluate the consistency between text inputs and the generated assets, we calculate the CLIP similarity between input text and rendered images. To better illustrate the results, we utilize three different CLIP models for testing, including ViT-L/14 (V-L/14), ViT-B/16 (V-B/16) and ViT-B/32 (V-B/32), respectively. The detailed results are displayed in Table 2. We compare representative text-to-3D approaches (e.g. WonderJ [55], Text2Room [16], and DreamFusion [39]) and Blender-driven 3D generation works (e.g. BlenderGPT, 3D-GPT [45], and SceneCraft [19]). Although the similarity scores of text-to-3D methods are higher, it is reasonable because their training or optimization includes the text-to-image alignment process. Compared to the Blender-driven 3D generation works, Scene achieves the highest score, indicating its capability to accurately execute the input prompts and generate results.
Efficiency Evaluation. To illustrate the efficiency of Scene, we provide the time required of our method compared with Infinigen [42] and human craft. The experiments are performed on a server equipped with dual Intel Xeon Processors (Skylake architecture), each with 20 cores, totaling 80 CPU cores. Additionally, we consult 3D PCG experts to determine the time needed to construct the same natural and urban scenes. The comparison results are shown in Table 3. From the results, we can observe that Scene is 7 times faster than Infinigen [42] in generating a scene with 150m 150m size. We also provide nearly 30 times faster than human experts creating a large-scale city by hand. This demonstrates the impressive efficiency of Scene in both large-scale natural scene and urban scene generation.
| 150m 150m | 500m 500m | 2.5km 2.5km | ||||
| Scene | City | Scene | City | Scene | City | |
| Human | 1h | 40min | 3h | 4h | - | >3w |
| Infinigen [42] | 14min | - | - | - | - | - |
| Scene (Ours) | 2min | 1.5min | 10min | 6min | - | 20h |
Personalized Editing Results. To demonstrate the capability of our method for personalized editing, we conduct experiments on 3D asset generation guided by users’ instructions. The results are shown in Fig. 10. It is evident that the changes in the edited text are closely related to the modifications in 3D assets. Scene demonstrates a versatile, highly controllable, and personalized editing ability by manipulating 3D assets from various perspectives. These results demonstrate that our method supports user-instruction editing, significantly reducing the difficulty of 3D asset generation and accelerating the industrial production process.
4.3 Ablation Study
In order to analyze the impact of various components within the systematic template on the system, we conduct an ablation study based on the tree plugin in PCGBench. During the experiment, we utilize the dataset from Section 4.1 for testing, maintaining a consistent format. We incrementally add or remove different parts of the systematic template, using ER@1 and SR@1 as metrics to observe the impact of various components on the system. The results are shown in Table 4. It is evident from the results that augmenting the Task, Document, Examples and Role components contributes to an increase in ER@1 and SR@1. Among these, The inclusion of the Example component results in the most significant improvement, resulting in a maximum increase of 76.00% and 51.09% in ER@1 and SR@1 respectively. Conversely, the Role component has the least impact, with maximum increases of 2.00% and 2.59% in ER@1 and SR@1 after its addition. These results suggest that the setting of Task significantly impacts the performance of LLMs. By clarifying the goals and requirements of the task, the system can better understand the user’s intent and better meet the needs when generating output. Examples play an important role in designing proxy prompts. By providing concrete examples, the system can better understand the user’s intent and needs and produce high-quality output related to the input text. The inclusion of documentation is also important for designing proxy prompts. Documentation provides background information that can help the system better meet user expectations when generating output.
| Task | Document | Examples | Role | ER@1 | SR@1 | |
| 1 | ✓ | 16.00 | 25.00 | |||
| 2 | ✓ | 42.00 | 47.62 | |||
| 3 | ✓ | ✓ | ✓ | 84.00 | 71.43 | |
| 4 | ✓ | 92.00 | 73.91 | |||
| 5 | ✓ | ✓ | 92.00 | 76.09 | ||
| 6 | ✓ | ✓ | 92.00 | 76.09 | ||
| 7 | ✓ | ✓ | ✓ | 92.00 | 78.26 | |
| 8 | ✓ | ✓ | ✓ | 94.00 | 78.72 | |
| 9 | ✓ | ✓ | ✓ | ✓ | 94.00 | 80.85 |
4.4 Comparing with Different LLMs
To investigate the performance of different variants of large language models (LLM) in Scene, we test public LLM APIs like gpt-3.5-turbo, gpt-4 and several external open-source LLMs in this subsection. To ensure the stability of LLM outputs, we set the temperature of LLM to 0 for all experiments. We conduct experiments on 3D scene and asset generation based on 50 scenario descriptions and 50 object descriptions in Sec. 4.1. The results are presented in Table 5. It is evident that gpt-4 delivers the most superior performance, with Mistral closely following as the second-best. Due to its superior performance and lower hardware requirements, the open-source Mistral is a highly appealing option. When compared to asset generation, the executability and success rates noticeably decline during the generation of large-scale natural scenes, a trend that can be attributed to the increased task complexity. In particular, as the number of components involved in the system expands, the LLM may face challenges in maintaining their accuracy. Nonetheless, our method exhibits consistent performance across different LLMs, maintaining high levels of executability and success rates.
| Model | Scene Generation | Asset Generation | ||
| ER@1 | SR@1 | ER@1 | SR@1 | |
| Llama2-7B[47] | 30.00 | 53.33 | 38.00 | 57.89 |
| Llama2-13B[47] | 44.00 | 59.09 | 54.00 | 66.66 |
| Mistral[24] | 76.00 | 68.42 | 94.00 | 85.11 |
| Gemma-2B[13] | 6.00 | 33.33 | 22.00 | 45.45 |
| Gemma-7B[13] | 36.00 | 55.56 | 68.00 | 73.53 |
| GPT-3.5-turbo[2] | 66.00 | 60.60 | 82.00 | 82.93 |
| GPT-4[37] | 86.00 | 86.05 | 96.00 | 85.42 |
5 Conclusion
In this paper, we introduce an advanced framework for large-scale scene generation, denoted as Scene. This framework automatically generates high-quality procedural models based on designers’ textual descriptions. The proposed method consists of two main components: PCGBench and PCGPlanner Agent. PCGBench includes an extensive repository of accessible procedural assets and thousands of meticulously crafted API documents. On the other hand, the PCGPlanner Agent can generate executable actions for Blender to create controllable and precise 3D assets following user instructions. Utilizing Scene, we can generate a city covering 2.5 km 2.5 km with accurate layout and geometric structures. This significantly reduces the time required, from two weeks for several professional PCG engineers to just a few hours for an ordinary user. Our method’s capabilities in controllable large-scale scene generation and editing, including tasks like asset placement and season translation, are validated through extensive experiments.
Limitations Although the proposed Scene is effective for the LLM-driven procedural large-scale scene generation, it still faces several limitations: 1) The performance of our method relies on the pretrained LLM model. This reliance may constrain the framework from generalizing to a wider range of applications. 2) We have collected a limited number of assets and APIs, which limits the variety of generated scenes and degree of action space. In the future, it is promising to fine-tune the LLMs with professional modeling knowledge and empower them to be a professional modeling assistant. Collecting more resources for PCGBench can also enhance the scene fidelity and aesthetics.
6 Acknowledgement
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0116500), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), and in part by the 2035 Innovation Program of CAS and the InnoHK program.
Appendix 0.A Supplementary Materials
In this paper, we propose an advanced Scene for large-scale scene generation according to the user’s instructions. The reproducible code is attached to the supplementary material, and some other important details are as follows.
0.A.1 PCGBench Collection
Blender Market offers a wide range of PCG plugins and 3D asset resources. To meet the demand for generating large-scale natural and urban scenes, we constructed PCGbench from four directions: Materials and Shading, Modeling, Rendering, and Asset Placement. Specifically, for generation-based PCG, we collected PCG models such as terrain, trees, flowers, rocks, and buildings. For material-based PCG, we gathered modifiable materials such as plant leaves, plant roots, terrain, buildings, snow, and sky. For asset placement-based PCG, we collected various placement PCG methods including slope-based, height-based, object-based, and geometric space-based placements. In terms of rendering-based PCG, we collected PCG resources related to rendering engines, rendering settings, and renderer optimizations. Additionally, for 3D asset resources, we collected a wide range of categories including buildings, vegetation, urban infrastructure, and textures sourced from the extensive realm of the internet.
Raw API collection: These PCG plugins are predominantly categorized into two types: one based on geometric node groups and the other based on Python code.It is necessary to gather API-related information from these PCG plugins, including API names, API function descriptions, API parameter descriptions, API parameter value ranges, and API parameter default values. For plugins based on geometric node groups, in order to obtain their original APIs, we need to test the geometric node groups of the plugins in Blender software and document their original APIs along with the corresponding information. For plugins based on Python code, we need to test the source code and record the functions within the plugin code as the original APIs, while also documenting the relevant information for these functions.
PCGBench: In order to optimize the usage of PCG plugins in PCGPlanner, we need to simplify each plugin. The simplification rules are as follows:
For generation-based plugins, we extract the APIs related to grid generation and remove irrelevant APIs. The criteria for successful extraction is the ability to independently generate the corresponding 3D grid using the plugin.
For material-based plugins, we extract the APIs that provide material variations and remove irrelevant APIs. The criteria for successful extraction is the ability to independently change the corresponding material using the plugin.
For placement-based plugins, we extract the APIs related to placement functionality and remove irrelevant APIs. Additionally, to accommodate multiple placement plugins, we have designed a universal collection format. The criteria for successful extraction is the ability to independently place objects in the current scene using the plugin.
For rendering-based plugins, we extract the APIs that provide scene rendering capabilities and remove irrelevant APIs. The criteria for successful extraction is the ability to independently perform the rendering tasks for the entire 3D scene using the plugin.
Furthermore, to better integrate and utilize all the PCG plugins and 3D assets, we actively utilize Blender’s API to create some basic functions such as importing, deleting, moving, and deforming. These APIs can be accessed and called by PCGPlanner to achieve better results. 3D Asset Dataset: In this study, we have developed a comprehensive dataset consisting of a large collection of 3D assets, including 1908 meticulously crafted models and 1294 textures. To facilitate text-to-text and text-to-image retrieval for the CLIP model, we have extensively annotated each 3D asset and supplemented them with rendered images.
0.A.2 Algorithm
In this chapter, we provide the pseudocode for the entire framework implementation workflow to clearly illustrate the logical relationship between the loops between the various agents. First, we use a loop to check the termination condition, i.e. stop the execution of the algorithm when = True. In each loop, we first call the dispatch agent, pass the input and Prompt to the agent, and receive a set of rough plan objects . Next, we iterate over each object and use the specialist agent to further break it down into a detailed plan to get a set of . We then execute each detailed plan and call Agent_retrieval and CLIP to search for the corresponding assets and APIs in PCGBench. Finally, we evoke execution Agent to manipulate each retrieved API , until all the sub-task is completed and change the to True.
0.A.3 Examples of Prompt Details
To provide a more intuitive demonstration of how agent prompts are designed and executed, we are presenting a series of examples. In Fig. 8 9 10 7, and 11, we show prompts for PCG plugins with different functionalities, along with the corresponding API calls for each PCG plugin. We also provide a tested example for each PCG component.
These images illustrate that our systematized prompt framework enables the successful execution of different tasks. Furthermore, the adoption of a unified prompt template for all agents simplifies maintenance and facilitates consistent invocation.
0.A.4 More Visualization Results
In Fig 12, we present the extended capabilities of Scene in urban and natural scene generation. The results demonstrate that our proposed Scene framework is capable of generating highly realistic scenes in both natural and urban environments. Furthermore, the generated content accurately corresponds to the provided texture descriptions. These outcomes serve as compelling evidence of the effectiveness and efficiency of our proposed LLM-based automatic 3D scene generation framework. In Fig 13, we show scenes generated by Scene without the utilization of any materials, highlighting the partial PCG components employed in the scene generation process, particularly the placement functionality achieved through geometric nodes. The results indicate that the generation of complex scenes requires the use of several sophisticated PCG plugins. Moreover, the generated scenes are solely based on object meshes, rather than relying on materials or rendering techniques. This implies that the scenes generated by Scene possess depth and other essential information, which holds significant implications across various domains.